OpenStudy (anonymous):

Please Help! Find all solutions in the interval [0, 2π). 4 sin2 x - 4 sin x + 1 = 0

OpenStudy (anonymous):

@zaynahf

OpenStudy (zehanz):

Do you perhaps mean:\[4\sin^{2}x-4sinx+1=0\]? In that case, you can treat it as an ordinary quadratic equation, with sinx as variable. If you like, you can temporarily set p=sinx. The equation then becomes:\[4p^2-4p+1=0\]Then solve for p. Later you can replace p with sinx again and solve for x.